# How can Diffusion Models Evolve into Continual Generators?

### v2

@@ Appendix C Model Architecture Diagram
Figure 4: Overall architecture of the Continual Consistency Diffusion (CCD) model. The backbone follows a U-Net-style encoder–decoder structure with residual and attention units, and is augmented with additional heads to support intermediate knowledge consistency objectives.
@@
Figure 4 presents a detailed schematic of the diffusion backbone used in all of our experiments. The design follows the “U-Net with cross-task hooks” blueprint popularised in contemporaneous diffusion work, but is augmented with three novel pathways that are required by the Continual Consistency Diffusion (CCD) training objective.
Appendix D Visualizations
@@
Figure 5: Comparison of generated results in 0-th task of MNIST.
@@
Figure 6: Comparison of generated results in the 0-th task of CIFAR100.
@@
Figure 7: Comparison of generated results in the 0-th task of Flowers102.
@@
Figure 8: Comparison of generated results in the 0-th task of CUB200.
@@
We conduct comprehensive evaluations across four continual generation benchmarks (MNIST, CIFAR100, Flowers102, and CUB200) to assess the effectiveness of our approach in retaining generative knowledge across tasks. On MNIST (Fig. 5), we observe that standard buffer-based baselines such as ER and A-GEM suffer from severe forgetting: they completely lose the ability to generate digits from the first task, including digits 0 and 1. In contrast, our method successfully reconstructs digit 0, evidencing improved knowledge retention. Nonetheless, the failure to accurately reproduce digit 1 suggests that catastrophic forgetting still persists, highlighting the need for more principled strategies for generative memory consolidation.
@@
On the more complex CIFAR100 dataset (Fig. 6), all compared methods, including ours, fail to retain generative knowledge from the initial task. This failure can be attributed to the minimal overlap in semantic content across tasks, making cross-task knowledge retention particularly challenging. These results underscore a key limitation of current approaches, including ours: the reliance on shared structural information for knowledge retention. In scenarios where such structure is absent, task interference remains severe. This raises an important open question, how can we effectively preserve and transfer independent, task-specific knowledge without impeding the acquisition of new information?
@@
In the Flowers102 benchmark (Fig. 7), where the dataset size is comparable to the buffer capacity, all methods achieve moderate generative performance. However, qualitative differences are evident. Our model consistently generates samples with higher visual fidelity and stronger alignment to real data. For instance, it successfully captures rare instances, such as white flowers in the third category, that A-GEM entirely fails to reproduce. Moreover, in categories prone to error (e.g., the seventh category), our model avoids semantic drift and maintains accurate class representation, suggesting a stronger capacity for handling underrepresented classes.
@@
Finally, on CUB200, a fine-grained benchmark (Fig. 8), our method clearly outperforms baselines in generative memory retention. It successfully reconstructs samples from the initial task, while ER and A-GEM fail to recover any meaningful representations. The alignment between fine-grained structure and our design principle of knowledge propagation yields consistently better generative fidelity. These findings not only validate our theoretical formulation but also demonstrate the practical advantage of our method in continual generation settings that demand nuanced representation learning.
@@
In summary, our approach shows strong resilience to forgetting, particularly in tasks with shared visual structure or fine-grained semantics. However, the limitations observed in unstructured task regimes, such as CIFAR100, emphasize the need for future work to address the preservation of task-specific knowledge in the absence of inter-task alignment.
Appendix E Future Improvements
(1) Hyper-parameter Sensitivity: The method depends on three hyper-parameters, whose optimal values vary across datasets. For instance, while MNIST settings may be effective, other datasets perform best with all values around $71e-$ 7. Future work should explore adaptive tuning strategies to improve robustness across datasets. (2) Buffer Quality: The metrics on CIFAR100-10T reveals that a larger buffer size (2560 or 5120) underperforms compared to a smaller buffer (512), suggesting that for certain datasets, the quality of stored samples is more critical than quantity.

### v3

@@ Appendix B Cross-Task Diffusion Evolution
Figure 6: Overall architecture of the CCD framework. The backbone follows a U-Net-style encoder–decoder structure with residual and attention units, and is augmented with additional heads to support intermediate knowledge consistency objectives.
@@ Appendix C Model Architecture Diagram
Figure 6 presents a detailed schematic of the diffusion backbone used in all of our experiments. The design follows the “U-Net with cross-task hooks” blueprint popularised in contemporaneous diffusion work, but is augmented with three novel pathways that are required by the Continual Consistency Diffusion (CCD) training objectives.
Appendix D Method-specific Implementation Details
Baseline Method Implementations
ER: Implements a FIFO-based buffer to store past samples, where the replay batch is combined with the current task batch to match the training batch size used in non-buffer-based methods, ensuring strict consistency. During each update, replayed samples are concatenated with the current batch for joint training.
Buffer Size (512 or 2560 or 5120):
Controls the number of stored past samples. Larger buffers reduce GCF but increase memory usage.
Replay Batch Size (100):
Number of samples drawn from memory per step.
L2 Regularization: Prevents GCF by constraining current model parameters to remain close to those from the previous task. The method maintains a frozen teacher model from the previous task and applies L2 penalty on parameter deviations: $\mathcal{L}_{L2}=\sum_{i}\|\theta_{i}-\theta_{teacher,i}\|_{2}^{2}$ with weight $\lambda_{L2}=50.0$. This direct parameter constraint helps preserve previous task knowledge while allowing adaptation to new tasks.
A-GEM: Uses gradient projection with an episodic memory to prevent interference with previous tasks. It computes a reference gradient on replay data and projects the current gradient to ensure a non-negative dot product with the reference.
Buffer Size (512 or 2560 or 5120):
Representative samples from prior tasks; more memory improves gradient accuracy.
Gradient Projection Rule:
Ensures updates do not decrease past-task performance, balancing stability and plasticity.
DCM: Adopts a hierarchical memory structure that organizes stored samples into adaptive clusters for diverse and efficient replay. It dynamically creates, merges, or updates clusters based on knowledge discrepancy measures.
Buffer Size (512 or 2560 or 5120):
Total number of stored samples across all clusters, defining overall replay capacity.
Cluster Capacity (64 or 128 or 256):
Maximum samples per cluster; when exceeded, the most redundant sample is removed.
Expansion Threshold (1500 or 2000 or 2500):
A new cluster is created if a sample is farther than this threshold from all current prototypes.
Maximum Clusters (20):
Upper bound on the total number of clusters; exceeding this triggers merging of the two most similar clusters.
Prototype Selection (Square_Error):
Each cluster maintains a prototype minimizing intra-cluster distances, updated periodically to reduce computation cost.
EWC: Estimates parameter importance via the Fisher Information Matrix; a quadratic penalty (diagonal approximation) constrains critical parameter drift.
Regularization Weight $\lambda_{EWC}=5.0$:
Balances old-task retention and new-task learning; higher values increase stability but reduce adaptability.
Fisher Diagonal:
Measures parameter sensitivity to past tasks, guiding which weights are most protected.
MAS: Measures parameter importance via the L2 norm of model outputs with respect to each parameter and accumulates importance across training.
Regularization Weight $\lambda_{MAS}=5e-5$:
Higher values preserve past knowledge more aggressively but may restrict new learning.
SI: Tracks parameter trajectories to compute importance scores, which are used to regularize updates after each task.
Regularization Weight $\lambda_{SI}=5.0$:
Controls the trade-off between stability and plasticity.
Stability Factor $\epsilon_{SI}=0.01$:
Prevents division by near-zero parameter changes, ensuring numerical stability.
LwF: Also known as Knowledge Distillation (KD), this method preserves knowledge from previous tasks by aligning the current model’s outputs with those of a frozen teacher from the prior task, using an MSE loss on noise predictions weighted by $\lambda_{LwF}=0.01$.
C-LoRA: A parameter-efficient continual learning approach that combines full fine-tuning for the first task with Low-Rank Adaptation for subsequent tasks. The method employs the following strategy:
Task 0:
Full fine-tuning of the entire UNet model to establish a strong backbone.
Subsequent Tasks:
Inject LoRA adapters with rank $r=8$, $\alpha=8$, dropout=0.1, targeting modules [’to_k’, ’to_q’, ’to_v’, ’to_out.0’, ’conv1’, ’conv2’].
Selective Training:
Only LoRA parameters and class embeddings are trainable; backbone remains frozen.
Memory Management:
Past class embeddings are preserved to prevent weight decay during training.
Task-specific Inference:
For each task, the model reconstructs the appropriate architecture by loading the backbone and injecting task-specific LoRA weights.
CCD Implementation Details
Our Continual Consistency Diffusion implements three consistency losses with specific computational strategies:
Inter-task Knowledge Consistency ($\mathcal{L}_{IKC}$): Employs Bregman divergence with Fisher Information-based preconditioner. The Fisher diagonal is computed as the squared mean of teacher model outputs, providing curvature-aware distance metrics between teacher and student score functions.
Unconditional Knowledge Consistency ($\mathcal{L}_{UKC}$): Applies time-dependent masking with threshold $t<700$ to emphasize semantically critical mid-diffusion phases. The reverse-time mean functions are computed using the reparameterization $\mu(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}x_{t}-\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}(1-\bar{\alpha}_{t})}}\epsilon_{\theta}(x_{t},t)$.
Prior Knowledge Consistency ($\mathcal{L}_{PKC}$): Utilizes KL divergence between mid-diffusion features at timestep $t=0$. The loss enforces semantic consistency by comparing prior embeddings from current and past models on replay samples.
Hyperparameter Configuration
Table 3 presents the hyperparameters used across all methods, extracted from our experimental codebase.
Parameter Value Description
CCD Framework
$\kappa$ (IKC weight) $1\times 10^{-5}\sim 1\times 10^{-7}$ Inter-task consistency
$\lambda$ (UKC weight) $1\times 10^{-5}\sim 1\times 10^{-7}$ Unconditional consistency
$\eta$ (PKC weight) $1\times 10^{-5}\sim 1\times 10^{-7}$ Prior knowledge consistency
Regularization Methods
EWC weight ($\lambda_{EWC}$) 5.0 Fisher penalty coefficient
L2 weight ($\lambda_{L2}$) 50.0 Parameter regularization
MAS weight ($\lambda_{MAS}$) $5\times 10^{-5}$ Importance-weighted penalty
SI weight ($\lambda_{SI}$) 5.0 Synaptic importance penalty
SI epsilon ($\epsilon_{SI}$) 0.01 Numerical stability term
LwF weight ($\lambda_{LwF}$) 0.01 Knowledge distillation penalty
DCM Configuration
Cluster capacity 64/128/256 Max samples per cluster
Expansion threshold 1500/2000/2500 New cluster creation threshold
Maximum clusters 20 Upper bound on cluster count
KDM type Square_Error Knowledge discrepancy measure
C-LoRA Configuration
LoRA rank ($r$) 8 Low-rank approximation rank
LoRA alpha ($\alpha$) 8 Scaling parameter
LoRA dropout 0.1 Dropout rate for LoRA layers
Memory & Training Setup
Buffer sizes 512/2560/5120 Rehearsal memory capacity
Replay batch size 100 Memory sampling size
Training batch size 200 Total training batch size
Diffusion timesteps 1000 Forward process steps
Inference steps 50 DDIM sampling steps
Learning rate $1\times 10^{-3}$ Adam optimizer
Table 3: Complete hyperparameter configuration for all implemented methods.
Appendix E Perceptual Metrics
In this section, we select three representative benchmarks to examine changes in perceptual loss, specifically the LPIPS metric. For LPIPS, we adapt the FID computation procedure to the CDG pipeline and compare against the most relevant baseline, ER. Detailed results are shown in Table. E.
From the results, we observe that LPIPS shows limited variation under our CCD framework compared to ER. This is largely attributed to inherent limitations of the UNet architecture and the pixel resolution used during training and evaluation. Notably, the baseline ER method already performs poorly on OxfordPets-5T and Flowers102-10T, posing additional challenges for CCD optimization. This highlights a key direction for future work, developing more capable model architectures to further improve perceptual quality.
[…]
MP↓ IMP↓ MP↓ IMP↓ MP↓ IMP↓
[…]
ER 0.38 0.54 0.79 0.78 0.65 0.68
[…]
ER 0.37 0.43 0.78 0.78 0.63 0.66
[…]
ER 0.37 0.42 0.78 0.77 0.64 0.65
[…]
Table 4: LPIPS comparison across datasets and buffer sizes.
Appendix F Visualizations
@@
Figure 7: Comparison of generated results in 0-th task of MNIST-5T.
@@
Figure 8: Comparison of generated results in the 0-th task of OxfordPets-5T.
@@
Figure 9: Comparison of generated results in the 0-th task of CIFAR100-10T.
@@
Figure 10: Comparison of generated results in the 0-th task of Flowers102-10T.
(a) The Original Results.
(b) Our CCD Results.
(c) ER Results.
(d) A-GEM Results.
Figure 11: Comparison of generated results in the 0-th task of CUB200-10T.
@@
We conduct comprehensive evaluations across five continual generation benchmarks (MNIST-5T, OxfordPets-5T, CIFAR100-10T, Flowers102-10T, and CUB200-10T) to assess the effectiveness of our approach in retaining generative knowledge across tasks. On MNIST-5T (Figure. 7), we observe that standard buffer-based baselines such as ER and A-GEM suffer from severe forgetting: they completely lose the ability to generate digits from the first task, including digits 0 and 1. In contrast, our method successfully reconstructs digit 0, evidencing improved knowledge retention. Nonetheless, the failure to accurately reproduce digit 1 suggests that GCF still persists, highlighting the need for more principled strategies for generative memory consolidation.
On the OxfordPets-5T dataset (Figure 8), our method demonstrates clear improvements over both ER and A-GEM. The samples produced by ER and A-GEM suffer from significant distortions and blurring, particularly evident in columns 2–5, where the cats’ faces often appear grotesquely warped and nearly unrecognizable. In contrast, our approach markedly reduces these artifacts, yielding substantially more realistic reconstructions. However, some residual imperfections in fine-grained details remain, suggesting that there is still considerable room for further enhancement.
@@
On the more complex CIFAR100-10T dataset (Figure. 9), all compared methods, including ours, fail to retain generative knowledge from the initial task. This failure can be attributed to the minimal overlap in semantic content across tasks, making cross-task knowledge retention challenging. These results underscore a key limitation of current approaches, including ours: the reliance on shared structural information for knowledge retention. In scenarios where such structure is absent, task interference remains severe. This raises an important open question, how can we effectively preserve and transfer independent, task-specific knowledge without impeding the acquisition of new information?
@@
In the Flowers102 benchmark (Figure. 10), where the dataset size is comparable to the buffer capacity, all methods achieve moderate generative performance. However, qualitative differences are evident. Our model consistently generates samples with higher visual fidelity and stronger alignment to real data. For instance, it successfully captures rare instances, such as white flowers in the third category, that A-GEM entirely fails to reproduce. Moreover, in categories prone to error (e.g., the seventh category), our model avoids semantic drift and maintains accurate class representation, suggesting a stronger capacity for handling underrepresented classes.
@@
Finally, on CUB200-10T, a fine-grained benchmark (Figure. 11), our method clearly outperforms baselines in generative memory retention. It successfully reconstructs samples from the initial task, while ER and A-GEM fail to recover any meaningful representations. The alignment between fine-grained structure and our design principle of knowledge propagation yields consistently better generative fidelity. These findings not only validate our theoretical formulation but also demonstrate the practical advantage of our method in continual generation that demand nuanced representation learning.
@@
In summary, our approach shows strong resilience to forgetting, particularly in tasks with shared visual structure or fine-grained semantics. However, its limitations in unstructured task regimes like CIFAR100-10T highlight the need for future work to better preserve task-specific knowledge in the absence of inter-task alignment.
Appendix G Future Improvements
While our CCD framework demonstrates significant advances in CDG pipeline, several avenues for improvement emerge from our theoretical analysis and experimental findings:
Adaptive Hyperparameter Tuning: Our method relies on three key hyperparameters ($\kappa$, $\lambda$, $\eta$) whose optimal values exhibit dataset-dependent variation. Future work should investigate meta-learning approaches or automated hyperparameter optimization strategies to enhance cross-dataset robustness and reduce manual tuning overhead.
Enhanced Buffer Construction: Our experiments reveal that buffer quality often supersedes quantity, as evidenced by CIFAR100-10T where smaller buffers (512) outperform larger ones (2560 or 5120). Although our proposed HDB shows promise on coarse-grained datasets, it exhibits limitations on fine-grained tasks due to the non-discriminative nature of intermediate diffusion representations. Future research should focus on developing more sophisticated sample selection mechanisms that better capture the semantic diversity essential for effective continual generation.
Minimal Cross-Task Overlap Scenarios: A fundamental limitation of our approach lies in scenarios with minimal semantic overlap between tasks, as demonstrated by the challenging CIFAR100-10T results. Our consistency-based framework inherently relies on shared knowledge structures, making it less effective when such commonalities are absent. Developing strategies for knowledge retention and propagation under conditions of minimal cross-task alignment represents a critical research direction.
Discriminative Representation Enhancement: The effectiveness of memory-based methods in diffusion models is constrained by the non-discriminative nature of intermediate representations, which consist primarily of isotropic Gaussian noise. Future work should explore techniques to enhance the discriminative quality of diffusion latent spaces, potentially through architectural modifications that preserve semantic information throughout the denoising process.